Please help with the second part of 8,# 9 and #10

A cube with a mass of 55 g is made from a metal wi
Answer #1

Volume = mass / density

V = 55 / 6 = 9.17 cm^3

V = L^3 = 9.17

L = 2.09 cm ..........Ans

9. Q = m C deltaT

= 380 x 1 cal/g deg C x (31 - 20)

= 4180 Cal


10. (A) Q = 90 x 1 x (75- 50) = 2250 cal

(B) Q = 90 x 1 x (50-25) = 2250 ca;

(c) remains unchangd.